### v1

# Gravitational-wave memory effects in the Damour-Esposito-Farèse extension of Brans-Dicke theory

Abstract
Gravitational-wave memory effects are lasting changes in the strain and its time integrals. They can be computed in asymptotically flat spacetimes using the conservation and evolution equations in the Bondi-Sachs framework. Modified theories of gravity have additional degrees of freedom with their own asymptotic evolution equations; these additional fields can produce differences in the memory effects in these theories from those in general relativity. In this work, we study a scalar-tensor theory of gravity known as the Damour-Esposito-Farèse extension of Brans-Dicke theory. We use the Bondi-Sachs framework to compute the field equations in Bondi-Sachs form, the asymptotically flat solutions, and the leading gravitational-wave memory effects. Although Damour-Esposito-Farèse theory has additional nonlinearities not present in Brans-Dicke theory, these nonlinearities are subleading effects; thus, the two theories share many similarities in the leading (and some subleading) solutions to hypersurface equations, asymptotic symmetries, and types of memory effects. The conservation equations for the mass and angular momentum aspects differ between the two theories, primarily because of the differences in the evolution equation for the scalar field. This leads to differences in the time dependence of the gravitational-wave memory signals that are produced during the quasicircular inspiral of compact binaries. These differences, however, are of second-order in a small coupling parameter of these theories, which suggests that it would be challenging to use memory effects to distinguish between these two theories.
I Introduction
Gravitational wave (GW) observations opened a new window to test general relativity (GR) in the strong-field regime of GR Abbott et al. (2019a, 2021a, 2021b). They are also allowing the nature of compact objects to be investigated (e.g., Isi et al. (2019); Bhagwat et al. (2020); Isi et al. (2021); Calderón Bustillo et al. (2021)), enabling the properties of the astrophysical population of merging black holes to be inferred Abbott et al. (2019b, 2021c, 2023a) and constraining the Hubble constant Abbott et al. (2017). These results were obtained from the first three observing runs of the LIGO-Virgo-KAGRA (LVK) collaboration, during which almost one-hundred compact-binary coalescences have been observed Abbott et al. (2019c, 2021d, 2023b).
During the fourth observing run, as the LVK detectors improve, the number of detections is expected to more than double Abbott et al. (2016). In the next decade, several new ground-based and space-based detectors are anticipated to operate (e.g., Amaro-Seoane et al. (2017); Punturo et al. (2010); Yagi (2013); Luo et al. (2016); Gong et al. (2021); Reitze et al. (2019)), which will both allow new sources to be detected in new frequency ranges and permit GWs from compact binaries to be detected at much larger rates and with significantly higher precision. Particularly the high-precision measurements will permit GR to be tested more precisely and to probe GW phenomena that have not yet been detected. The currently unobserved phenomena that will be the focus of this paper are GW memory effects.
What is now referred to as the displacement GW memory effect was first studied in the 1970s Zel’dovich and Polnarev (1974) (see also Smarr (1977); Turner (1978); Epstein (1978); Bontz and Price (1979) and earlier related discussion in Newman and Penrose (1966)). The study of GW memory effects has expanded since their first computation several decades ago largely because of developments in their theoretical study and observational search prospects.
The observational prospects for detecting the displacement memory had been studied prior to the observation of GWs Braginsky and Grishchuk (1985); Braginsky and Thorne (1987); Kennefick (1994); Favata (2009a); Pollney and Reisswig (2011), but these studies focused on individual sources of GWs. After the detection of binary-black-hole (BBH) mergers by LIGO, it was realized that detecting evidence for the memory effect in the entire population of BBH mergers was possible Lasky et al. (2016). Subsequent forecasts based on the BBH population inferred from the LVK detections have indicated that the memory effect in the BBH population is likely to be observed during the fifth LVK observing run Hübner et al. (2020); Boersma et al. (2020); Grant and Nichols (2023). There are active searches for the memory effect in BBH populations using LVK data Hübner et al. (2020, 2021); Cheung et al. (2024) and from individual events from subsolar mass BBHs Ebersold and Tiwari (2020) or with pulsar timing arrays Wang et al. (2015); Agazie et al. (2024). It will also be possible to detect the memory with future GW detectors on the ground Johnson et al. (2019); Grant and Nichols (2023); Goncharov et al. (2024) and in space Favata (2009a); Islo et al. (2019); Goncharov et al. (2024); Gasparotto et al. (2023); Inchauspé et al. (2024).
There have been developments in the study of memory effects from a theoretical perspective, too. First, GW memory effects were found to be part of an infrared triangle (see, e.g., Strominger and Zhiboedov (2016); Strominger (2017)) which relates them to the asymptotic symmetries of the spacetime and to soft theorems of gravitational scattering Weinberg (1965). Second, several additional GW memory observables were predicted that generalize the displacement memory. The displacement memory is named as such because it can be measured by the change in separation of two comoving freely falling observers after the passage of a burst of GWs (i.e., their geodesic deviation). If the observers have a relative velocity, then there will be a residual change in their displacement after the burst of waves that depends on their initial velocity and the time integral of the shear. This subleading GW memory Flanagan et al. (2019); Grant and Nichols (2022) (also called drift memory Grant (2024)) is closely connected with the spin Pasterski et al. (2016); Flanagan and Nichols (2017); Nichols (2017) and center-of-mass (CM) Nichols (2018) GW memory effects. Observers with relative acceleration can measure a hierarchy of higher GW memory effects, which depend on the initial time derivatives of the relative acceleration and two or more integrals of the GW strain Flanagan et al. (2019); Grant and Nichols (2022); Grant (2024); Grant and Mitman (2024); Siddhant et al. (2024).
All of these GW memory effects can be computed from (i) “fluxes” (or “pseudofluxes”), which are nonlinear in the Bondi metric functions and vanish in the absence of radiation, and (ii) “charges,” which can be constructed to be constant in the absence of radiation. In some cases, the charges and fluxes are associated with underlying symmetries of asymptotically flat spacetimes. This is shown most definitively for the displacement memory effect, which is closely connected to the supertranslation symmetries of asymptotically flat spacetime in the Bondi-Metzner-Sachs (BMS) group Bondi et al. (1962); Sachs (1962) via the infrared triangle discussed above. The spin and CM memory effects are related to the extended Barnich and Troessaert (2010a, b, 2011) or generalized Campiglia and Laddha (2014, 2015) BMS algebras, though not quite as straightforwardly as with the standard BMS group Compère et al. (2018). The charges from which the higher memory effects also may be related to a $w_{1+\infty}$ algebraic structure Strominger (2021); Freidel et al. (2022, 2024); Compère et al. (2022); Geiller (2024). However, in this paper we will restrict to computations of the displacement and spin memory effects sourced by the nonlinear Christodoulou (1991); Blanchet and Damour (1992); Flanagan and Nichols (2017); Nichols (2017) (or flux) terms.
We focus on the leading displacement and spin memory effects in this paper, because our focus will be on memory effects in gravitational theories beyond GR. GWs from compact-binary sources have been studied extensively in these theories (see, e.g., Yunes and Siemens (2013); Yunes et al. (2024)), and they have been constrained by the LVK observations Abbott et al. (2021a, b). Memory effects in beyond-GR theories, however, have been studied in less detail, but there have been a handful of calculations performed in the past few years. Brans-Dicke (BD) theory Tahura et al. (2020, 2021); Hou and Zhu (2020a, b) and dynamical Chern-Simons gravity Hou et al. (2022a, b) were studied using the Bondi-Sachs framework. The displacement GW memory effect also was computed in most general scalar-vector-tensor theory Heisenberg et al. (2023, 2024), which itself encompasses a wide class of specific beyond-GR theories, using the Isaacson effective stress-energy tensor in beyond-GR theories, instead. There have also been linearized-theory calculations of asymptotic symmetries and their connections to GW memory effects in the Lorentz-violating Einstein-Æther theory Hou et al. (2024) and in diffeomorphism-invariant theories containing tensor degrees of freedom that closely resemble those in GR Hou (2024).
In both GR and beyond GR theories, computing memory effects can be performed perturbatively, approximately using balance laws, or numerically in full generality. In GR, memory effects have been computed during the inspiral phase of a binary system using the flux terms and post-Newtonian (PN) Favata (2009b, 2011); Ebersold et al. (2019) or self-force Elhashash and Nichols (2024); Cunningham et al. (2024) waveforms. These calculations are consistent with full PN calculation at 4PN order that do not make use of the balance laws Trestini and Blanchet (2023); Blanchet et al. (2023a, b) when compared at the overlapping PN orders for each calculation. Post-Newtonian calculations using the balance laws have also been performed in BD theory Tahura et al. (2021). The fluxes in the Bondi-Sachs framework also can be used to compute memory effects approximately using full, numerical-relativity (NR) waveforms Nichols (2017); Mitman et al. (2021, 2020); Khera et al. (2021); Grant and Mitman (2024), because NR simulations that extract waveforms using extrapolation often do not capture GW memory effects (see, e.g., Boyle et al. (2019)).
Although NR simulations using extraction do not capture the displacement memory effect, there are two closely related numerical frameworks that do. The first is Cauchy-characteristic evolution (CCE), which evolves NR data with a finite outer boundary to future null infinity after the finite-domain NR simulation has been run, and Cauchy-characteristic matching (CCM), which performs both evolutions coupled and simultaneously (see, e.g., the review Winicour (2009)). CCE and CCM compute gravitational waveforms at future null infinity that include memory effects in a particular BMS frame (see the review Mitman et al. (2024) for more details about the choice of the frame). The NR code SpECTRE Moxon et al. (2020, 2023); Ma et al. (2024a); Deppe et al. (2024), for example, is under development and is implementing both the CCE and CCM algorithms.
While most of the development of CCE and CCM has been performed in the context of GR, there has been recent work in Ma et al. (2024b), which solved the Einstein-Klein-Gordon system (by incorporating a massless scalar field into the SpECTRE CCE). This will be an important step towards computing GW memory effects from compact binary coalescences in non-GR theories, which often have additional scalar degrees of freedom. More specifically, the work of Ma et al. (2024b) could be applied to Bergmann–Wagoner scalar-tensor (ST) theories Wagoner (1970); Bergmann (1968) or Damour-Esposito-Farèse (DEF) theory with a single scalar field Damour and Esposito-Farese (1992), in the Einstein frame. NR simulations of black-hole–neutron-star systems in DEF theory were performed, which could be used as input to CCE (see, e.g., Ma et al. (2023)).
Useful cross-checks of the GW memory effects computed from these anticipated CCE and CCM simulations in ST theories will be having expressions for the conservation equations that the memory effects must satisfy and having analytical PN results of GW memory effects. Our focus in this paper will be to derive such balance laws and compute the PN memory waveforms in DEF theory with a single scalar field in Bondi-Sachs coordinates, which are used by the CCE and CCM frameworks. We will focus on the displacement and spin memory effects, as they are the leading electric and magnetic-parity memory effects that have been computed in GR (see, e.g., Nichols (2017); Mitman et al. (2021); Grant and Mitman (2024); Siddhant et al. (2024)) as well as in BD theory Tahura et al. (2020, 2021).
DEF theory is a useful example of a ST theory to consider because it contains additional nonlinearities that produce nontrivial strong-field dynamics (such as spontaneous scalarization for neutron stars Damour and Esposito-Farese (1993), and it also reduces to BD theory when the coupling function in the action between gravity and the scalar field becomes a constant. The Bondi-Sachs framework is well adapted for studying nonlinear dynamics of massless fields in the wave zone, and it will allow us to understand what strong-field features of DEF theory also have an impact on the propagation of tensor and scalar GWs far from the source. Because it reduces to BD theory in a particular limit, we can also more easily determine which features of the memory waveforms are associated with having an additional, noninteracting scalar degree of freedom, and which are associated with the additional nonlinearities in the dynamics of the scalar field.
In this paper, we will focus on the similarities and differences of aspects of asymptotically flat spacetimes and memory effects in DEF and BD theories. We will write the results in a way that highlights the similarities with the results of Refs. Tahura et al. (2020, 2021) when they occur and makes manifest the differences between the two theories. Most of these differences arise from the evolution of the scalar field and, in particular, the subleading dynamical part of the field. We will also discuss how these differences ultimately enter into the PN inspiral memory waveforms that are computed from compact-binary sources on quasicircular orbits.
Specifically, the organization of this paper is as follows: In Sec. II, we review aspects of ST theories in the Einstein frame in Bondi-Sachs coordinates. We use the Einstein-frame results to determine appropriate asymptotically flat boundary conditions and the corresponding form of the metric in the Jordan frame. We then compute the field equations of DEF theory in the Jordan frame and derive asymptotically flat solutions. In Sec. III, we discuss the leading and subleading tensor and scalar memory effects in DEF theory. We then describe how the tensor effects can be computed using the conservation equations for the mass and angular momentum aspects. The scalar effects are computed from charge contributions, and the subleading scalar has a new nonlinear contribution in DEF theory that we discuss. We also describe how the PN inspiral memory signals from compact binaries differ between DEF and BD theories. We present our conclusions in Sec. IV. Finally, we discuss BMS symmetries and the transformation of the Bondi metric functions and scalar fields under these symmetries in Appendix A.
Throughout this paper, we use units in which the speed of light is $c=1$ and in which the gravitational constant in the Einstein frame satisfies $\tilde{G}_{E}=1$. We use the conventions for the metric and curvature tensors given in Misner et al. (1973). Greek indices ($\mu,\nu,\alpha,\dots$) represent four-dimensional spacetime indices, and uppercase Latin indices $(A,B,C,\dots)$ represent indices on the 2-sphere. We used the xAct suite Martín-García (2025, 2008) to perform some of the calculations. A Mathematica notebook with these calculations is available in this GitHub repository Tahura (2025).
II Bondi-Sachs framework for DEF theory
In this section, we review the formulation of DEF theory and give the forms of the field equations in Bondi-Sachs coordinates. We will typically refer to the theory as DEF theory, though we may also refer to it as ST theory, as is done in Ma et al. (2023). First, we give a brief overview of DEF theory in the Einstein frame in Sec. II.1; the form of the field equations and the assumptions that we make for the boundary conditions for asymptotic flatness are identical to those in BD theory Tahura et al. (2020); Hou and Zhu (2020a). Next, we transform to the Jordan frame to obtain the corresponding fall-off conditions of the Bondi metric functions and the scalar field; the transformation is of precisely the same form as in BD theory Tahura et al. (2020). Finally, we compute the field equations in the Jordan frame and derive asymptotically flat solutions in Sec. II.3.
II.1 Einstein-frame field equations and fall-off conditions
In the Einstein frame, the action of a ST theory in the absence of additional matter fields takes the same form as in BD theory Wagoner (1970); Berti et al. (2015):
$S=\int d^{4}x\sqrt{-\tilde{g}}\left[\frac{\tilde{R}}{16\pi}-\frac{1}{2}\tilde{g}^{\rho\sigma}\left(\tilde{\nabla}_{\rho}\Phi\right)\left(\tilde{\nabla}_{\sigma}\Phi\right)\right].$ (1)
We use the notation $\tilde{g}_{\mu\nu}$ for the metric in the Einstein frame, $\tilde{R}$ for the Ricci scalar, and ~$\tilde{\nabla}_{\mu}$ for the covariant derivative compatible with $\tilde{g}_{\mu\nu}$. The field $\Phi$ is a massless real scalar, and we set the “bare” gravitational constant [^1] The phrase “bare” gravitational constant refers to the constant that appears in the action. The locally measured value of Newton’s constant can differ from the bare one in modified theories of gravity. $\tilde{G}_{E}$ in the Einstein frame equal to one. Varying the action with respect to the metric and the scalar field gives the field equations of DEF theory:
$\tilde{G}_{\mu\nu}=8\pi\tilde{T}_{\mu\nu}^{(\Phi)},\qquad\tilde{\nabla}_{\mu}\tilde{\nabla}^{\mu}\Phi=0.$ (2a)
Here $\tilde{G}_{\mu\nu}=\tilde{R}_{\mu\nu}-(1/2)\tilde{R}\,\tilde{g}_{\mu\nu}$ is the Einstein tensor, and the stress-energy tensor $\tilde{T}_{\mu\nu}^{(\Phi)}$ has the usual form for that of a massless, real scalar field $\Phi$:
$\tilde{T}_{\mu\nu}^{(\Phi)}=\tilde{\nabla}_{\mu}\Phi\tilde{\nabla}_{\nu}\Phi-\tilde{g}_{\mu\nu}\left[\frac{1}{2}\tilde{g}^{\alpha\beta}\tilde{\nabla}_{\alpha}\Phi\tilde{\nabla}_{\beta}\Phi\right].$ (2b)
We now introduce Bondi coordinates $(\tilde{u},\,\tilde{r},\,\tilde{x}^{A})$, which are well adapted for studying outgoing radiation from an isolated source. The variable $\tilde{u}$ is the retarded time, $\tilde{r}$ is a radial coordinate, and $\tilde{x}^{A}$ are coordinates on two-sphere surfaces of constant $\tilde{u}$ and $\tilde{r}$ ($A$ can be either $1$ or $2$). The explicit form of the metric is (see Bondi et al. (1962); Sachs (1962); Mädler and Winicour (2016))
$ds^{2}=$ $-\frac{\tilde{V}}{\tilde{r}}e^{2\tilde{\beta}}d\tilde{u}^{2}-2e^{2\tilde{\beta}}d\tilde{u}d\tilde{r}$
$+\tilde{r}^{2}\tilde{h}_{AB}\left(d\tilde{x}^{A}-\tilde{U}^{A}d\tilde{u}\right)\left(d\tilde{x}^{B}-\tilde{U}^{B}d\tilde{u}\right).$ (3)
The four Bondi metric functions—$\tilde{V}$, $\tilde{\beta}$, $\tilde{U}^{A}$ and $\tilde{h}_{AB}$, which depend on all four Bondi coordinates—contain six degrees of freedom, because the symmetric tensor $\tilde{h}_{AB}$ contains only two degrees of freedom when the determinant condition of Bondi gauge is imposed: namely,
$\det\left[\tilde{h}_{AB}\right]=q\left(\tilde{x}^{C}\right).$ (4)
The function $q\left(\tilde{x}^{C}\right)$ is commonly selected to be the determinant of the round two-sphere metric $q_{AB}$, which gives $\tilde{r}$ the interpretation of being an areal radius. The form of the metric in Eq. (II.1) already has the remaining three independent Bondi gauge conditions imposed:
$\tilde{g}_{\tilde{r}A}=\tilde{g}_{\tilde{r}\tilde{r}}=0.$ (5)
These conditions ensure that the hypersurfaces of constant $\tilde{u}$ are null and that the angular coordinates $\tilde{x}^{A}$ are constant along outgoing null rays.
Note that the field equations in Eq. (2) have the same form as the Einstein-Klein-Gordon system in GR with a massless real scalar field, precisely as in BD theory. We thus assume the same fall-off for $\Phi$ as a power series expansion in $1/\tilde{r}$, with the leading-order piece being constant, as in Tahura et al. (2020):
$\Phi\left(\tilde{u},\tilde{r},\tilde{x}^{A}\right)=\Phi_{0}+\frac{\Phi_{1}\left(\tilde{u},\tilde{x}^{A}\right)}{\tilde{r}}+O(\tilde{r}^{-2}).$ (6)
We also impose the same fall-off conditions on metric functions as in Tahura et al. (2020):
$\tilde{h}_{AB}={}$ $q_{AB}(\tilde{x}^{C})+\frac{\tilde{c}_{AB}(\tilde{u},\,\tilde{x}^{C})}{\tilde{r}}+\mathcal{O}\left(\frac{1}{\tilde{r}^{2}}\right),$ (7a)
$\tilde{\beta}={}$ $\frac{\tilde{\beta}_{1}(\tilde{u},\,\tilde{x}^{C})}{\tilde{r}}+\mathcal{O}\left(\frac{1}{\tilde{r}^{2}}\right)\,,$ (7b)
$\tilde{U}^{A}={}$ $\frac{\tilde{U}_{2}^{A}(\tilde{u},\,\tilde{x}^{C})}{\tilde{r}^{2}}+\mathcal{O}\left(\frac{\log\tilde{r}}{\tilde{r}^{3}}\right),\,$ (7c)
$\tilde{V}={}$ $\tilde{r}+\tilde{V}_{0}(\tilde{u},\,\tilde{x}^{C})+\mathcal{O}\left(\frac{1}{\tilde{r}}\right)\,.$ (7d)
The coefficient of the $1/\tilde{r}$ piece of conformal 2-metric, $\tilde{c}_{AB}$, is the shear tensor, which is traceless with respect to $q_{AB}$ (i.e., $q^{AB}\tilde{c}_{AB}=0$); this follows from the determinant condition of Bondi gauge. The shear tensor encodes the two transverse-traceless degrees of freedom of tensor GWs.
Because the fall-off conditions and the evolution equations have precisely the same form as in BD theory, the solutions to the field equations will also have the same form. For example, they give that $\tilde{\beta}_{1}=0$ and $\tilde{U}_{2}^{A}=-(1/2)\eth_{B}\tilde{c}^{AB}$, where ð$\eth_{B}$ is covariant derivative compatible with $q_{AB}$. A more detailed discussion is given in Tahura et al. (2020). As in Tahura et al. (2020), the detailed form of the equations is not essential for the remaining calculations, because we use the Einstein-frame results to determine the fall-off conditions in the Jordan frame. We will solve the field equations in the Jordan frame rather than transforming the full solution in the Einstein frame to the Jordan frame.
II.2 Conformal transformation to the Jordan frame and Jordan-frame fall-off conditions
The spacetime metric in the Einstein and Jordan frames are related by a conformal transformation Damour and Esposito-Farese (1992); Carroll (2019):
$\tilde{g}_{\mu\nu}=\lambda g_{\mu\nu}.$ (8)
The function $\lambda(x^{\mu})$ is the scalar field in the Jordan frame, and in generic scalar-tensor theories, it is related to $\Phi$ through a differential equation involving a “coupling” function $\omega(\lambda)$:
$\frac{d\Phi}{d\log\lambda}=\sqrt{\frac{3+2\omega(\lambda)}{16\pi}}.$ (9)
By postulating the relationships in Eqs. (8) and (9), the Einstein-frame action in Eq. (1) can be transformed to the action in the Jordan frame,
$S=\frac{1}{16\pi}\int d^{4}x\sqrt{-g}\left[\lambda R-\frac{\omega(\lambda)}{\lambda}g^{\mu\nu}\nabla_{\mu}\lambda\nabla_{\nu}\lambda\right]$ (10)
(see, e.g., Damour and Esposito-Farese (1992)). Here $\nabla_{\mu}$ is the covariant derivative compatible with the Jordan frame metric $g_{\mu\nu}$ and $R$ is the corresponding Ricci scalar. Given that we chose $\tilde{G}_{E}=1$ in the Einstein frame, the scalar field is dimensionless (rather than having units of the inverse of Newton’s constant, as in BD theory Brans and Dicke (1961)).
In BD theory, $\omega(\lambda)$ is assumed to be a constant, which allows Eq. (9) to be integrated and $\log\lambda$ is proportional to $\Phi$ minus its asymptotic value. In DEF theory, a similar relationship between $\log\lambda$ and $\Phi$ is postulated, but instead of assuming that $\log\lambda$ is linear in $\Phi$, the Einstein-frame scalar $\Phi$ is Taylor expanded to quadratic order around its asymptotic value $\Phi_{0}$ Damour and Esposito-Farese (1993). We use the notation of Ma et al. (2023) to write the expression as
$\log\lambda=-4\sqrt{\pi}\alpha_{0}(\Phi-\Phi_{0})-4\pi\mathcal{B}_{0}(\Phi-\Phi_{0})^{2},$ (11)
where $\alpha_{0}$ and $\mathcal{B}_{0}$ are the two coupling constants of DEF theory. By computing $d\log\lambda/d\Phi$ from Eq. (11), the inverse can be substituted into the differential equation (9) and solved for $\omega(\lambda)$, with $\Phi(\lambda)$ being given implicitly (see Ma et al. (2023)):
$\omega(\lambda)=-\frac{3}{2}+\frac{1}{2\left[\alpha_{0}+2\sqrt{\pi}\mathcal{B}_{0}(\Phi-\Phi_{0})\right]^{2}}\,.$ (12)
Note that DEF theory reduces to BD theory when $\mathcal{B}_{0}$ goes to zero and $\alpha_{0}=-(3+\omega_{\text{BD}})^{-1/2}$, and that it reduces to GR when $\mathcal{B}_{0}=\alpha_{0}=0$.
We now use the expansion of the Einstein-frame scalar field and metric functions, the conformal transformation of the metric, and the relationship between $\lambda$ and $\Phi$ in Eq. (11) to determine the asymptotic boundary conditions on the fields $\lambda$ and $g_{\mu\nu}$ in the Jordan frame. Substituting Eq. (6) into (11) and exponentiating gives
$\lambda=1-\frac{4\sqrt{\pi}\alpha_{0}\Phi_{1}}{\tilde{r}}+\mathcal{O}\left(\tilde{r}^{-2}\right).$ (13)
Given that the differences between DEF and BD theories arise at quadratic order in the scalar field, it is reasonable that the scalar field has the same asymptotic behavior as in BD theory: namely, it goes as a constant plus a function of the coordinates $\tilde{u}$ and $\tilde{x}^{A}$ normalized by $\tilde{r}$. Note, however, that the leading order piece of $\lambda$ has been chosen to be one here, whereas in Tahura et al. (2020) it was allowed to be an arbitrary constant $\lambda_{0}$. Thus, to compare results here with those in Tahura et al. (2020), one should set $\lambda_{0}=1$ in the results of Tahura et al. (2020). [^2] In ST theories, the value of $\lambda_{0}$ is related to the asymptotic value of the gravitational constant, $G$ by $G=(4+2\omega_{0})/[\lambda_{0}(3+2\omega_{0})]$. To set $G=1$ as $r\rightarrow\infty$, then $\lambda_{0}$ would be different from one, in general (see, e.g., the discussion after Eq. (4.6) of Ref. Mirshekari and Will (2013) and also footnote 5 of Ref. Tahura et al. (2020)). Note, however, that the GR limit is $\omega_{0}\rightarrow\infty$, which leads to $G=1$ with $\lambda_{0}=1$. In this paper, we set $\lambda_{0}=1$ because we do not require the asymptotic value of $G$ to be 1 at infinity. Thus, the asymptotic boundary conditions on the fields in the Jordan frame in DEF theory will be the same as those in BD theory, which are described in more detail in Tahura et al. (2020).
We now remind the reader that when performing the conformal transformation in Eq. (8) to obtain the Jordan-frame metric, the Bondi gauge conditions $g_{\tilde{r}\tilde{r}}=g_{\tilde{r}A}=0$ are satisfied, but the determinant condition is not satisfied: namely the determinant $\det[g_{AB}]=\tilde{r}^{4}\lambda^{-2}q(x^{B})$ is no longer independent of $\tilde{u}$ and its $\tilde{r}$ dependence is no longer an overall multiplicative factor of $\tilde{r}^{4}$. To restore the determinant condition, we use the same transformation used in BD theory in Tahura et al. (2020), where we define new coordinates by
$u=\tilde{u},\quad r=\frac{\tilde{r}}{\sqrt{\lambda}},\quad x^{A}=\tilde{x}^{A}.$ (14)
In the new set of coordinates, the Jordan frame metric can be written in Bondi-Sachs form as
$g_{\mu\nu}dx^{\mu}dx^{\nu}=$ $-\frac{V}{r}e^{2\beta}du^{2}-2e^{2\beta}dudr$
$+r^{2}h_{AB}\left(dx^{A}-U^{A}du\right)\left(dx^{B}-U^{B}du\right),$ (15)
where $\det[h_{AB}]=q(x^{C})$. Using the conformal transformation in Eq. (8), the coordinate transformation in Eq. (14), and the boundary conditions on the metric and scalar field in the Einstein frame [given in Eqs. (7) and (6)], we find the Jordan frame metric functions and scalar field should have the following scaling with $r$:
$\lambda={}$ $1+\frac{\phi}{r}+\mathcal{O}\left(\frac{1}{r^{2}}\right),$ (16a)
$h_{AB}={}$ $q_{AB}+\frac{c_{AB}}{r}+\mathcal{O}\left(\frac{1}{r^{2}}\right),$ (16b)
$\beta={}$ $-\frac{1}{2r}\phi+\mathcal{O}\left(\frac{1}{r^{2}}\right)\,,$ (16c)
$U^{A}={}$ $\frac{1}{2r^{2}}(\eth^{A}\phi-\eth_{B}c^{AB})+\mathcal{O}\left(\frac{\log r}{r^{3}}\right),$ (16d)
$V={}$ $(1+\partial_{u}\phi)r+\mathcal{O}(r^{0})\,.$ (16e)
Two-sphere indices were raised with $q^{AB}$ in the expressions above. The two-sphere tensor $c_{AB}$ is trace-free with respect to $q_{AB}$ (i.e., $q^{AB}c_{AB}=0$), and it has the same form as in the Einstein frame [i.e., $c_{AB}(u,x^{A})=\tilde{c}_{AB}(\tilde{u},\tilde{x}^{A})$]. We have also defined $\phi=-4\sqrt{\pi}\alpha_{0}\Phi_{1}$.
Because we will solve the Jordan-frame field equations in the next section, the emphasis here is not on determining the detailed relationship between the Einstein-frame and Jordan-frame metric quantities and scalar fields (the leading-order results are the same as in BD theory; see Tahura et al. (2020)). Rather, we are just determining what are the expected fall off rates in $r$ of the metric functions in the Jordan frame. For example, by enforcing the determinant condition of the Bondi gauge in the coordinates ($u,\,r,\,x^{A}$), the leading order part of the metric component $g_{uu}$ becomes a function of $u$ and $x^{A}$ (this would not occur if we used a gauge that did not enforce this condition; see Hou and Zhu (2020a)). Because scalar-tensor theories admit a scalar “breathing-mode” polarization of GWs, the time dependence of the leading part of $g_{uu}$ is where this radiative degree of freedom enters the metric if the Bondi determinant condition is enforced; the curvature tensors and other properties of the solution are consistent with our notions of asymptotically flat boundary conditions defined in the earlier parts of this section.
II.3 Jordan-frame field equations
In this section, we solve the field equations in Jordan frame using the Bondi coordinates and the boundary conditions on the fields that were discussed in Sec. II.2. We first give the field equations in the Jordan frame, and then we determine solutions to these equations in which the fields are expanded as a series in $1/r$.
II.3.1 Field equations in Bondi gauge
The modified Einstein equations are obtained from varying the Jordan-frame action in Eq. (10) with respect to the inverse metric. The scalar field equation is obtained from varying the action with respect to $\lambda$ (up to boundary terms) and combining the result with the trace of the modified Einstein equations. The resulting field equations are of the form
$G_{\mu\nu}$ $=\frac{1}{\lambda}\left(8\pi T_{\mu\nu}^{(\lambda)}+\nabla_{\mu}\nabla_{\nu}\lambda-g_{\mu\nu}\Box\lambda\right),$ (17a)
$\Box\lambda$ $=-\frac{1}{3+2\omega(\lambda)}\frac{d\omega}{d\lambda}\nabla_{\mu}\lambda\nabla^{\mu}\lambda.$ (17b)
Here $G_{\mu\nu}=R_{\mu\nu}-(1/2)R\,g_{\mu\nu}$ is the Einstein tensor, and $\Box=g^{\mu\nu}\nabla_{\mu}\nabla_{\nu}$ is the wave operator constructed using the Jordan frame metric. The stress-energy tensor of $\lambda$ is given by
$T_{\mu\nu}^{(\lambda)}=\frac{\omega(\lambda)}{8\pi\lambda}\left(\nabla_{\mu}\lambda\nabla_{\nu}\lambda-\frac{1}{2}g_{\mu\nu}\nabla^{\alpha}\lambda\nabla_{\alpha}\lambda\right).$ (18)
The modified Einstein equations (17a) have the same form as the ones in BD theory, except the factor of $\omega$ is now $\lambda$-dependent. The scalar field equation in BD theory is $\Box\lambda=0$, which is consistent with the fact that $\omega$ is independent of $\lambda$ in BD theory. In DEF theory, there is a nonzero term on the right side, which depends on $\lambda$. It will be at least quadratic in $\lambda$ and its derivatives.
To help in referring to specific components of the modified Einstein equations, we introduce the notation
$\mathcal{G}_{\mu\nu}\equiv G_{\mu\nu}-\frac{1}{\lambda}\left(8\pi T_{\mu\nu}^{(\lambda)}-\nabla_{\mu}\nabla_{\nu}\lambda-g_{\mu\nu}\Box\lambda\right)=0.$ (19)
For the subsequent calculations, it will also be useful to write (12) explicitly as a function of $\log\lambda$ rather than implicitly, as it is currently written. To do so, we solve the quadratic equation for $\Phi-\Phi_{0}$ given in Eq. (11). The solution consistent with our assumptions for $\Phi$ and $\lambda$ in Eqs. (6) and (11) is given by
$\Phi-\Phi_{0}=\frac{\alpha_{0}}{2\sqrt{\pi}\mathcal{B}_{0}}\left(-1+\sqrt{1-\frac{\mathcal{B}_{0}}{\alpha_{0}{}^{2}}\log\lambda}\right).$ (20)
Substituting this expression into Eq. (12) gives the result for $\omega$ as a function of $\lambda$. Given that $\lambda$ has an expansion of the form in Eq. (13), then $\log\lambda$ will be a small parameter. We can then Taylor expand the result in Eq. (20) in $\log\lambda$; at the order in the expansion in $1/r$ in which we work in the remainder of the paper, it will be sufficient to work to linear order in $\log\lambda$. Thus, we have that
$\omega(\lambda)=-\frac{3}{2}+\frac{1}{2\alpha_{0}^{2}}+\frac{\mathcal{B}_{0}}{2\alpha_{0}{}^{4}}\log\lambda+\mathcal{O}[(\log\lambda)^{2}].$ (21)
Because the modified Einstein equations (17a) in DEF theory have the same form as in BD theory when the constant $\omega$ of BD theory is replaced by the function $\omega(\lambda)$, the field equations formally have the same form under this replacement. The explicit form of the field equations in the Jordan frame were given in the Appendix of Tahura et al. (2020). Because there are some subtleties related to integrating the “hypersurface equations” that were not discussed there, we describe this in more detail below.
First, the $\mathcal{G}_{rr}=0$ components of the field equations give rise to a hypersurface-type equation
$2\left(\frac{2}{r}+\partial_{r}\log\lambda\right)\partial_{r}\beta=$
$\frac{1}{\lambda}\partial^{2}_{r}\lambda+\frac{\omega}{\lambda^{2}}(\partial_{r}\lambda)^{2}+\frac{1}{4}h^{AB}h^{CD}\partial_{r}h_{AC}\partial_{r}h_{BD}.$ (22)
For general solutions (those that are not necessarily asymptotically flat), the term multiplying $\partial_{r}\beta$ could be zero, which prevents one from generically dividing by this term and integrating $\partial_{r}\beta$ to obtain $\beta$ as a first integral of $\lambda$, $h_{AB}$ and their radial derivatives on a surface of fixed retarded time $u$. However, for the asymptotically flat solutions that we consider in this paper, the term $\partial_{r}\log\lambda$ will scale as $1/r^{2}$, so the multiplicative factor will be nonzero for finite $r>0$, and Eq. (II.3.1) can be integrated as a hypersurface equation as in GR or in DEF theory in the Einstein frame.
Next, the $\mathcal{G}_{rA}=0$ components of the field equations are those that produce a hypersurface equation for $U^{A}$. The form of this equation is given by
$\frac{1}{2r^{2}}\left(\partial_{r}+\partial_{r}\log\lambda\right)(r^{4}e^{-2\beta}h_{AB}\partial_{r}U^{B})=$
$r^{2}\partial_{r}\left(\frac{1}{r^{2}}D_{A}\beta\right)-\frac{1}{2}h^{BC}D_{B}\left(\partial_{r}h_{AC}\right)-\frac{D_{A}\lambda}{\lambda r}$
$+\frac{\omega}{\lambda^{2}}\partial_{r}\lambda D_{A}\lambda-\frac{1}{\lambda}D_{A}\beta\partial_{r}\lambda-\frac{1}{2\lambda}h^{BC}D_{B}\lambda\partial_{r}h_{AC}$
$+\frac{1}{\lambda}\partial_{r}D_{A}\lambda.$ (23)
Here $D_{A}$ is the covariant derivative compatible with metric $h_{AB}$. Note that unlike in GR or in DEF theory in the Einstein frame in which the only term in the differential operator acting on $r^{4}e^{-2\beta}h_{AB}\partial_{r}U^{B}$ is the $\partial_{r}$ term, there is now a second term involving $\partial_{r}\log\lambda$ on the left-hand side. Thus, in general, to solve for $\partial_{r}U^{B}$ (and subsequently $U^{B}$) one would need to invert the more complicated differential operator that acts on $r^{4}e^{-2\beta}h_{AB}\partial_{r}U^{B}$ (instead of just integrating with respect to $r$ as in GR or DEF theory in the Einstein frame). However, for the asymptotically flat boundary conditions that we consider in this paper, the second term involving $\partial_{r}\log\lambda$ will be subleading to the first term; thus, the hierarchy for solving for a given order in $1/r$ of the expansion of $U^{B}$ will be determined by the expansions of $h_{AB}$, $\lambda$, $\beta$, and lower order parts of $U^{B}$. This is a slight modification of the typical description of the hierarchy of the hypersurface equations.
The last of the hypersurface equations is the one for $V$, which can be obtained from the trace of the components of the field equations $\mathcal{G}_{AB}=0$. As with the previous hypersurface equations, the differential operator acting on $V$ will now be $\lambda$-dependent (unlike in GR or in DEF theory in the Einstein frame, but similarly to BD theory in the Jordan frame). In addition, the wave operator acting on the scalar field $\Box\lambda$ enters into the hypersurface equation for $V$. In BD theory, assuming the scalar-field equation of motion is satisfied in vacuum ($\Box\lambda=0$) then this term vanishes; in DEF theory, one can instead replace $\Box\lambda$ with the right-hand side of Eq. (17b) when the scalar-field equation is satisfied. Thus, the hypersurface equation for $V$ can be written in the form
$2\left[\partial_{r}+\partial_{r}\log\lambda\left(1-\frac{1}{4}\frac{\partial\log(2\omega+3)}{\partial\log r}\right)\right]V={}$ $e^{2\beta}[\mathscr{R}-2h^{AB}\left(D_{A}D_{B}\beta+D_{A}\beta D_{B}\beta\right)]+\frac{1}{r^{2}}D_{A}\partial_{r}(r^{4}U^{A})$
$-\frac{1}{2}r^{4}e^{-2\beta}h_{AB}\partial_{r}U^{A}\partial_{r}U^{B}-\frac{e^{2\beta}h^{AB}}{\lambda^{2}}(\omega D_{A}\lambda D_{B}\lambda+\lambda D_{B}D_{A}\lambda)$
$+\frac{r}{\lambda}\left[2(\partial_{u}\lambda+U^{A}D_{A}\lambda)+r(D_{A}U^{A})\partial_{r}\lambda\right]$
$-\frac{d\omega/d\lambda}{\lambda(2\omega+3)}[2r^{2}\partial_{u}\lambda\partial_{r}\lambda+2r^{2}\partial_{r}\lambda U^{A}D_{A}\lambda-e^{2\beta}h^{AB}D_{A}\lambda D_{B}\lambda].$ (24)
We introduced $\mathscr{R}$ to be the Ricci scalar of the metric $h_{AB}$. We also used the relation $\partial_{r}\omega(\lambda)=(\partial_{r}\lambda)d\omega/d\lambda$. In addition to the more complicated radial differential operator that needs to be inverted, note that the retarded time derivative, $\partial_{u}\lambda$, also appears in the hypersurface equation for $V$. This occurs both in BD theory and DEF theory in the Jordan frame (see the first term in the third line). The hypersurface equations in GR, or in the Einstein frame for DEF theory, can be integrated using the metric functions $\beta$, $U^{A}$ and $h_{AB}$ (and the scalar field $\lambda$ for DEF theory) and spatial derivatives of these fields. In both BD and DEF theories information about how the scalar field changes off of a hypersurface of constant $u$ is required to integrate the hypersurface equations. In the Jordan frame, one would generically need to simultaneously solve the hypersurface equation with the evolution equation for the scalar field. Thus, we give this evolution equation next.
To write the scalar field equation, it is helpful to note that in Grant and Nichols (2022), it was observed that the evolution equation for the trace-free part of $h_{AB}$, rather than being interpreted as an evolution equation, could also be considered to be a hypersurface equation for the trace-free part of $\partial_{u}h_{AB}$. In this spirit, and given that $\partial_{u}\lambda$ appears in the hypersurface equations, it is convenient to write the scalar-field evolution equation in this form, as follows:
$2r^{2}\left[\partial_{r}+\frac{1}{r}\left(1+\frac{1}{2}\frac{\partial\log(2\omega+3)}{\partial\log r}\right)\right]\partial_{u}\lambda=$ $-2rU^{A}(D_{A}\lambda+rD_{A}\partial_{r}\lambda)+e^{2\beta}h^{AB}(2D_{A}\beta D_{B}\lambda+D_{B}D_{A}\lambda)$
$-r^{2}D_{A}\lambda(\partial_{r}U^{A})-r^{2}D_{A}U^{A}\partial_{r}\lambda+(V+r\partial_{r}V)\partial_{r}\lambda+rV\partial_{r}^{2}\lambda$
$-\frac{d\omega/d\lambda}{(3+2\omega)}\left[-r(\partial_{r}\lambda)^{2}V-e^{2\beta}h^{AB}D_{A}\lambda D_{B}\lambda+2r^{2}(U^{A}D_{A}\lambda)\partial_{r}\lambda\right].$ (25)
Writing the scalar field equation in this form highlights the similar structure it has with the hypersurface equation for $V$ in Eq. (II.3.1). Without the assumption of asymptotic flatness, Eqs. (II.3.1) and (II.3.1) would need to be solved as two coupled hypersurface equations for $V$ and $\partial_{u}\lambda$.
With the asymptotically flat boundary conditions assumed in Sec. II.2, it will be possible to solve for the leading-order in $1/r$ evolution of the scalar field without needing $V$; then $V$ can be solved in terms of the leading part of $\partial_{u}\lambda$. This process can be iterated to obtain higher orders in $1/r$ of $\partial_{u}\lambda$ and then higher orders in $1/r$ of $V$. We describe the asymptotic solution of these hypersurface equations in more detail in the next part.
We do not give the evolution equations (the trace-free part of $\mathcal{G}_{AB}=0$) or the supplementary equations ($\mathcal{G}_{uu}=0$ and $\mathcal{G}_{uA}=0$) in full generality here. The expressions are lengthy, and in the solution with asymptotically flat boundary conditions, we will work at an order in $1/r$ in which there are no differences from the results in BD theory given in Tahura et al. (2020).
II.3.2 Asymptotically flat solutions in the Jordan frame
Given the structure of the hypersurface equations described above, if we postulate a form for the expansion of the 2-metric $h_{AB}$ and $\lambda$, then we can solve the hypersurface equations to determine corresponding expansions of $\beta$, $U^{A}$, $\partial_{u}\lambda$ and $V$. The evolution equations and supplementary equations give information about the dynamics of the metric functions and the scalar field.
Based on the fall-off conditions posed in the Einstein frame, which we transformed to the Jordan frame, we write, as in BD theory, the expansion of the 2-metric $h_{AB}$ and scalar field $\lambda$ as
$h_{AB}=q_{AB}+\frac{c_{AB}(u,x^{A})}{r}+\frac{d_{AB}(u,x^{A})}{r^{2}}+\mathcal{O}(r^{-3}),$
$\lambda(u,r,x^{A})=1+\frac{\lambda_{1}\left(u,x^{A}\right)}{r}+\frac{\lambda_{2}\left(u,x^{A}\right)}{r^{2}}+\mathcal{O}(r^{-3})\,.$
The determinant condition of Bondi gauge requires that $q^{AB}c_{AB}=0$. It also constrains the $d_{AB}$ trace portion (with respect to $q_{AB}$), but it does not constrain the trace-free part. We thus write $d_{AB}$ as
$d_{AB}=D_{AB}+\frac{1}{4}c_{CD}c^{CD}q_{AB},$ (27)
with $q^{AB}D_{AB}=0$.
Because $\omega$ is a function of $\lambda$, then the expansion in Eq. (26b) allows us to write $\omega$ as a series in $\lambda$. Substituting Eq. (26b) into Eq. (21) gives
$\omega=\omega_{0}{}+\frac{\omega_{1}{}}{r}+\mathcal{O}(r^{-2})$ (28)
with the coefficients in the expansion being
$\omega_{0}=\frac{1}{2}\left(-3+\frac{1}{\alpha_{0}^{2}}\right),\qquad\omega_{1}=\frac{\mathcal{B}_{0}}{2\alpha_{0}{}^{4}}\lambda_{1}.$ (29)
The fact that $\omega$ in Eq. (28) at leading order is a constant and differs from a constant at order $1/r$ allows us to estimate the orders in $1/r$ at which asymptotically flat solutions of DEF theory will first differ from those of BD theory.
Specifically, the right-hand side of Eq. (17b) for the evolution of the scalar field will be of order $1/r^{3}$. This implies that the evolution equation for $\lambda_{1}$ will be the same in BD and DEF theories, but the evolution of $\lambda_{2}$ will differ. [^3] In more detail, $\partial_{u}\lambda_{1}$ will be unconstrained “radiative data” in both BD and DEF theories. Having $\lambda_{1}$ being the free radiative data in both theories, however, does not imply that the value of $\lambda_{1}$ produced by “equivalent” sources would be the same. As we will discuss in Sec. III.3, for example, the tensor and scalar gravitational waveforms from compact binaries with the same numerical values of the masses and sensitivities in BD and DEF theories will differ. The fact that $\lambda_{2}$ differs from BD theory then allows us to determine the orders in $1/r$ for which there are differences in the stress-energy tensor $T_{\mu\nu}^{(\lambda)}$ and in the remaining terms on the right-hand side of the modified Einstein equation (17a). The leading-order terms will come from those linear in $\lambda$ in Eq. (17a), specifically from the $\lambda_{2}$ terms which evolve differently in BD theory from in DEF theory. Given its scaling with $1/r$, we anticipate differences to occur at order $1/r^{2}$ in $\beta$, $1/r^{3}$ in $U^{A}$, and $1/r^{2}$ in $h_{AB}$; they would also be expected to affect the conservation equations for the mass aspect and angular momentum aspect, which will be introduced later. [^4] These are just expectations based on scaling arguments, but there could be cancellations that arise that make the differences first enter at higher orders. For example, there will be no additional changes to the $1/r^{3}$ part of $U^{A}$ from solving the hypersurface equation in Eq. (II.3.1), because several terms proportional to the angular derivatives of $\lambda_{2}$ cancel; however $U^{A}$ will differ at order $1/r^{3}$ because the evolution of the angular momentum aspect does not have such a cancellation.
Given that we will need the evolution equation for $h_{AB}$ at leading and next-to-leading order in $1/r$, where we expect no difference from BD theory, it will be simplest to review the solutions of this equation first (although, as discussed in Tahura et al. (2020), for example, the hypersurface equations generally should be solved before the evolution equations are). As in BD theory, the leading-order in $1/r$ part of the evolution equation for $h_{AB}$ gives that $\partial_{u}c_{AB}=N_{AB}$, where $N_{AB}$ is the news tensor. The news (for short) is the free tensor radiative data in this formalism (i.e., the evolution equations do not constrain the evolution of the shear). The next order in $1/r$ of the evolution equation for $h_{AB}$ gives
$\partial_{u}\left(D_{AB}+\frac{1}{2}\lambda_{1}c_{AB}\right)=0.$ (30)
This implies that $D_{AB}+(\lambda_{1}/2)c_{AB}$ is a constant of motion (i.e., it is determined by the initial values of these fields). As we discuss below, we will chose the initial data so as to eliminate an order $r^{-3}\log r$ term in $U^{A}$.
In the discussion of the hypersurface equations, it was observed that, in general, solving the scalar-field evolution equation requires solving the $\beta$ and $U^{A}$ hypersurface equations before, and solving the hypersurface equation for $V$ simultaneously. However, given the expected fall-offs of the Bondi metric functions and the scalar field in Eq. (16), then as in BD theory, the leading-order part of the scalar-field equation shows that $\partial_{u}\lambda_{1}$ is unconstrained by the evolution equation (making it the free radiative data, which was referred to as the “scalar news” in Tahura et al. (2020)). Thus, $\lambda_{1}$ does not depend on the solutions to the hypersurface equations. The evolution of $\lambda_{2}$ will, so we postpone the discussion of its solution until after we solve the hypersurface equations.
Solving the hypersurface equations proceeds similarly to the procedure in BD theory. We can obtain the solution for $\beta$ expanded in a series in $1/r$ by substituting Eq. (26) into Eq. (II.3.1) and integrating $\partial_{r}\beta$ with respect to $r$ to obtain $\beta$. The solution for $U^{A}$ is obtained by using the result for $\beta$ and the expressions in Eq. (26) in the hypersurface equation (II.3.1). The resulting equation can be integrated twice with respect to $r$ to obtain a series expansion for $U^{A}$. This double integration permits the introduction of a function of integration at order $1/r^{3}$, which is denoted $L_{A}(u,x^{B})$ and called the angular momentum aspect. It also allows for a logarithm term at order $r^{-3}\log r$. However, as in BD theory, the coefficient of this term turns out to be non-dynamical. Specifically, it is the two-sphere divergence of the constant of motion in Eq. (30). For simplicity in Tahura et al. (2020), we chose initial data that admits expansions of the metric functions in $1/r$ (as in Bondi et al. (1962)) rather than the “polyhomogeneous” expansion with terms of the from $(\log r)^{m}/r^{n}$ (for $n>m$; see, e.g., Friedrich (1998)). We will also make this assumption of a smooth expansion in $1/r$ here. Finally, using $U^{A}$, $\beta$, $\partial_{u}\lambda_{1}$ and Eq. (26), we can substitute these expressions into the hypersurface equation for $V$ in Eq. (II.3.1) to obtain the leading and next-to-leading parts of $V$. Because only $\partial_{u}\lambda_{1}$ is required at these orders, the hypersurface equation has effectively decoupled from the scalar field equation. It is now possible to integrate $V$, which has a new function of integration, the mass aspect $M(u,x^{A})$. In summary, we find that the hypersurface equations have solutions